OpenStudy (anonymous):

Solve the equation √(5x + 10) + 8 = x. If there is only one solution, say so. Otherwise find the sum of the solutions.

OpenStudy (anonymous):

\[\sqrt(5x+10) = x - 8\] \[\sqrt(5x + 10) ^ 2 = (x - 8) ^2\] \[5x + 10 = x^2 - 16x + 64\] \[0 = x^2 - 21x + 54\] \[0 = (x-18)(x-3)\] Therefore x = 18 and x = 3 are possible solution. Check each solution. checking x = 18 \[\sqrt(5*18 + 10) = 18 - 8\] \[\sqrt(100) = 10\] \[10 = 10\] x = 18 is a solution checking x = 3 \[\sqrt(5*3+10) = 3 - 8\] square roots can't be negative. Therefore x = 3 is not a souliton

OpenStudy (anonymous):

Thanks so much, you're a lifesaver!